OpenStudy (anonymous):

find expression for apparent nth term of geometric sequence 2,-4,6,-8,10,... geometric sequence is first number times common ratio (ie. -2). negative common ratio for alternating negative-postive sequence. But more is needed.

OpenStudy (anonymous):

No way is this a geometric sequence.

OpenStudy (anonymous):

I assumed that. arithmetic sequence?

OpenStudy (anonymous):

not even arithmetic ... maybe harmonic??

OpenStudy (anonymous):

how to find the expression for that sequence?

OpenStudy (anonymous):

consider 2+4+6+8+10+12+... first term=2... c.d=4-2=2 \[tn=\left( -1 \right)^{n-1}\left\{ 2+\left( n-1 \right)2 \right\}=2\left( -1 \right)^{n-1}n\]

OpenStudy (anonymous):

correction i have taken series it is a sequence

OpenStudy (anonymous):

surjithayer, the sequence is 2,-4,6,-8,10

OpenStudy (anonymous):

insert commas inplace of +

OpenStudy (anonymous):

normally the expression would be first number * common ratio ^(n - 1) alternating pos/neg indicates negative common ratio but not in this case. I want to learn the steps if possible

OpenStudy (anonymous):

\[\left( -1 \right)^02,\left( -1 \right)^14,\left( -1 \right)^26,\left( -1 \right)^38,\left( -1 \right)^410,....\]

OpenStudy (anonymous):

interesting approach, but expression won't find nth term of sequence

OpenStudy (anonymous):

\[2\left\{ \left( -1 \right)^{1-1}1,\left( -1 \right)^{2-1}2,\left( -1 \right)^{3-1}3,\left( -1 \right)^{4-1}4,\left( -1 \right)^{5-1}5,...,\left( -1 \right)^{n-1}n,... \right\}\] \[tn=2\left( -1 \right)^{n-1}n\]

OpenStudy (anonymous):

thank you!! but how can I derive that formula myself? The usual geometric formula is ar^(n-1)

OpenStudy (anonymous):

no problem for 2,4,6,8,10,... for powers of -1,terms are 0,1,2,3,4,.... a=0,d=1-0=1 tn=a+(n-1)d=0+(n-1)=n-1 so power is n-1

OpenStudy (anonymous):

you got it.

OpenStudy (anonymous):

thanks so much!!

OpenStudy (anonymous):

yw